# Conjugacy properties of multivariate unified skew–elliptical distributions

### v2

@@ 2.1 Multivariate unified skew–elliptical distributions
${\bf{z}}\stackrel{{\scriptstyle d}}{{=}}{\bm{\xi}}+\bm{\omega}{\bf{z}}^{\star},\quad{\bf{z}}^{\star}\stackrel{{\scriptstyle d}}{{=}}(\tilde{\bf{z}}\mid\tilde{\bf{z}}_{0}+\bm{\tau}>\bm{0}),\qquad\mbox{with}\quad\begin{bmatrix}\tilde{\bf{z}}\\ \tilde{\bf{z}}_{0}\end{bmatrix}\sim\mbox{{EC}}_{m+q}\left(\begin{bmatrix}\bm{0}\\ \bm{0}\end{bmatrix},\begin{bmatrix}\bar{\bm{\Omega}}&{\bm{\Delta}}\\ {\bm{\Delta}}^{\top}&\bar{\bm{\Gamma}}\end{bmatrix},g^{(m+q)}\right).$ (5)
Such a representation is the one adopted by [5] for the sub–family of sun distributions and is particularly convenient for deriving the mean vector and covariance matrix of ${\bf{z}}$. More specifically, leveraging (5), the law of total expectation, and the previously–discussed closure properties of elliptical distributions, we have that
@@
where $\psi$ is a scalar whose specific form can be obtained from the derivations in Arellano-Valle and Genton [10]. The above expressions clarify that moments of sue random vectors can be directly obtained from those of multivariate truncated elliptical distributions [e.g., 13, 54, 39, 53, 59]. In addition, Equation (7) shows that to enforce a lack of correlation among the entries in ${\bf{z}}$, it is not sufficient to impose suitable diagonal or block–diagonal structures within $\bm{\Omega}$. Rather, these constraints should be combined with additional ones, e.g., on the shape matrix ${\bm{\Delta}}$. Such a result is useful in Section 3 to derive examples of practically–meaningful priors and likelihoods inducing sue posterior distributions. To this end, Sections 3.1, 3.2 and 3.3 state general conjugacy properties and then specialize these results to the two most widely–implemented examples of sue distributions that are presented in detail in Section 2.2 below, namely, multivariate unified skew–normals (sun) [5] and multivariate unified skew–$t$ (sut) [62]. The results for sun clarify that the conjugacy properties derived by [32, 36, 4] can be obtained as a special case, and under a different proof technique, of the more general sue framework introduced in the present article. Conversely, the conjugacy results stated for sut are a novel contribution that extends to a broader class of models of potential practical interest the findings in Song and Xia [57] and Zhang et al. [65] on specific Student’s $t$ linear regressions and multivariate probit formulations based on skew–elliptical link functions, respectively.
2.2 Relevant sub–classes of multivariate unified skew–elliptical distributions
@@
The sun and sut families arise from general skewed perturbations of multivariate Gaussians and Student’s $t$ densities, respectively. As such, these formulations are arguably the most relevant and practically–impactful sub–classes within the sue representation. In addition, recalling Arellano-Valle and Azzalini [5], Arellano-Valle and Genton [10], and Wang et al. [62], both sun and sut admit additive stochastic representations which allow for i.i.d. sampling under posterior distributions belonging to these two sub–classes, thus facilitating Bayesian inference; see also Yin and Balakrishnan [63] for a recent extension of these stochastic representations to more general multivariate skew–elliptical distributions, beyond sun and sut. Sections 2.2.12.2.2 provide a concise overview of sun and sut sub–classes, respectively. A more extensive treatment can be found in, e.g., [5] and [62].
2.2.1 Multivariate unified skew–normal distributions
@@
The sun family has been introduced by Arellano-Valle and Azzalini [5] to provide a single class of distributions capable of unifying several extensions of the original multivariate skew–normal [19]. Relevant examples of representations that belong to such a wide class are extended multivariate skew–normals [15, 14] and closed skew–normals [42, 43], among others. Recalling Arellano-Valle and Azzalini [5], a random vector ${\bf{z}}\in\mathbb{R}^{m}$ has multivariate unified skew–normal (sun) distribution, i.e., ${\bf{z}}\sim{\mbox{SUN}}_{m,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}})$, if its density $\mathcal{p}({\bf{z}})$ can be expressed as
@@
Although the above representation is originally derived in Arellano-Valle and Azzalini [5] under a selection representation similar to (5) applied to an underlying Gaussian random vector, the density in (8) can also be derived directly from (1) under a suitable choice of the density generators. In particular, let $g^{(m)}=\phi^{(m)}(u)=(2\pi)^{-m/2}\exp(-u/2)\mathds{1}_{u>0}$, $g^{(q)}=\phi^{(q)}(u)=(2\pi)^{-q/2}\exp(-u/2)\mathds{1}_{u>0}$ and recall that, in the particular Gaussian setting, the conditional generator $\phi^{(q)}_{Q({\bf{y}})}$ coincides with the unconditional one $\phi^{(q)}$. Then, recalling related derivations in Arellano-Valle and Genton [10], and replacing these density generators in (1), directly yields expression (8), thus clarifying that suns are special cases of sue distributions.
Recalling the discussion within Section 1, the above sun class has been at the basis of important recent advancements in conjugate Bayesian inference for a broad class of routinely–implemented representations relying on fully–observed, discretized or partially–discretized Gaussian and multivariate skew–normal models [32, 36, 4]. As discussed in Section 3, these conjugacy properties can be recovered as a special case of those we derive for the general sue family, which, in turn, allows us to extend such results to larger classes, including the sut one introduced in Section 2.2.2.
2.2.2 Multivariate unified skew–$t$ distributions
The success of the sun family has motivated several extensions aimed at deriving similar skewed representations for other sub–classes of the elliptical family. A noticeable and natural generalization is provided by the class of sut distributions [e.g., 62] which can be obtained by replacing Gaussians with suitably–defined multivariate Student’s $t$ within the original selection representation of suns. Recalling, e.g., Wang et al. [62], this yields a density for the multivariate unified skew–$t$ (sut) random vector ${\bf{z}}\sim{\mbox{SUT}}_{m,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},\nu)$, defined as
@@ 2.2.2 Multivariate unified skew–$t$ distributions
with $\alpha_{\nu,Q({\bf{z}})}=[\nu+Q({\bf{z}})]/{(\nu+m)}$, $Q({\bf{z}})=({\bf{z}}-{\bm{\xi}})^{\top}\bm{\Omega}^{-1}({\bf{z}}-{\bm{\xi}})$, and $\nu>0$ denoting the degrees of freedom. The remaining parameters have, likewise, similar roles and related interpretations to those in (8). Analogously, $t_{m}({\bf{z}}-{\bm{\xi}};\bm{\Omega},\nu)$, $T_{q}[\smash{\alpha_{\nu,Q({\bf{z}})}^{-1/2}}\{\bm{\tau}+{\bm{\Delta}}^{\top}\bar{\bm{\Omega}}^{-1}\bm{\omega}^{-1}({\bf{z}}-{\bm{\xi}})\};\bar{\bm{\Gamma}}-{\bm{\Delta}}^{\top}\bar{\bm{\Omega}}^{-1}{\bm{\Delta}},\nu+m]$ and $T_{q}(\bm{\tau};\bar{\bm{\Gamma}},\nu)$ correspond to the density and cumulative distribution functions, computed at ${\bf{z}}-{\bm{\xi}}$, $\smash{\alpha_{\nu,Q({\bf{z}})}^{-1/2}}\{\bm{\tau}+{\bm{\Delta}}^{\top}\bar{\bm{\Omega}}^{-1}\bm{\omega}^{-1}({\bf{z}}-{\bm{\xi}})\}$ and $\bm{\tau}$, respectively, of the centered multivariate Student’s $t$ distributions with scale matrices $\bm{\Omega}\in\mathbb{R}^{m\times m}$, $\bar{\bm{\Gamma}}-{\bm{\Delta}}^{\top}\bar{\bm{\Omega}}^{-1}{\bm{\Delta}}\in\mathbb{R}^{q\times q}$ and $\bar{\bm{\Gamma}}\in\mathbb{R}^{q\times q}$, and degrees of freedom $\nu$, $\nu+m$ and $\nu$, respectively. Notice that, for $q=1$, one retrieves the multivariate extended skew-$t$ in [9]. Moreover, when $\bm{\tau}=\bm{0}$ and ${\bm{\Delta}}=\bm{0}$, the numerator and denominator in (9) coincide, and, hence, the density reduces to that of a multivariate Student’s $t$ distribution $\mathcal{T}_{m}({\bm{\xi}},\bm{\Omega},\nu)$ with location ${\bm{\xi}}$, scale $\bm{\Omega}$, and degrees of freedom $\nu$. This implies that the multivariate Student’s $t$ is obtained as a special case of sut.
@@
As for the sun, also the sut density in Equation (9) can be derived from (1) under a suitable choice of the density generators. In particular, let $g^{(m)}=t_{v}^{(m)}(u)=c(\nu,m)[1+u/\nu]^{-(\nu+m)/2}$, and $g^{(q)}=t_{v}^{(q)}(u)=c(\nu,q)[1+u/\nu]^{-(\nu+q)/2}$, where the generic quantity $c(a,b)$ is defined as $c(a,b)=\Gamma[(a+b)/2]/[\Gamma(a/2)(\pi a)^{b/2}]$, and $\Gamma(\cdot)$ is the usual gamma function. Recalling Arellano-Valle and Genton [10], under these settings, the induced conditional density generator is equal to $t_{\nu,Q({\bf{z}})}^{(q)}(u)=\alpha_{\nu,Q({\bf{z}})}^{-q/2}t_{\nu+m}^{(q)}(\alpha_{\nu,Q({\bf{z}})}^{-1/2}u)$, where $\alpha_{\nu,Q({\bf{z}})}$ is defined as $\alpha_{\nu,Q({\bf{z}})}=[\nu+Q({\bf{z}})]/(\nu+m)$, whereas $Q({\bf{z}})$ denotes a quadratic form induced by the $m$–dimensional conditioning variable. Hence, replacing the generic density generators in (1) with those defined above yields, as a result of straightforward calculations, the sut density in Equation (9). Interestingly, as clarified in e.g., Wang et al. [62], when $\nu\rightarrow\infty$, this density reduces to (8), thereby establishing a direct connection between sut and sun distributions. This suggests that the conjugacy properties derived in [32, 36, 4] for sun might extend to sut, and, more generally, to sue distributions. Two promising results in this direction have been derived in Song and Xia [57] and in Zhang et al. [65], but only with a focus on Student’s $t$ linear regression and on specific multivariate binary data settings. Leveraging the sue properties in Section 2.3, we prove in Section 3 that sue conjugacy holds for a substantially larger class of models which further embraces formulations of potential interest in practice.
2.3 Properties of unified skew–elliptical distributions
@@
Lemmas 15 below state several central properties of sue distributions that are at the core of the novel conjugacy results derived in Section 3. More specifically, Lemmas 12 establish closure under linear combinations, marginalization, and conditioning of sue distributions. These properties have appeared also in Arellano-Valle and Genton [10], but under a different parameterization. Conversely, Lemmas 35 state novel results that are useful to study sue conjugacy under broad classes of models and to derive special cases of potential interest in practical contexts.
@@
Let ${\bf{z}}\sim{{\mbox{SUE}}}_{m,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},g^{(m+q)})$. In addition, denote with ${\bf{A}}\in\mathbb{R}^{r\times m}$ a matrix with rank $r\leq m$, and let $\bm{b}\in\mathbb{R}^{m}$ be a vector of constants. Then
@@
where ${\bm{\Delta}}_{\bf{A}}=\bm{\omega}_{\bf{A}}^{-1}{\bf{A}}\bm{\omega}{\bm{\Delta}}$ with $\bm{\omega}_{\bf{A}}=\mbox{diag}({\bf{A}}\bm{\Omega}{\bf{A}}^{\top})^{1/2}$. Moreover, let ${\bf{z}}_{C}\in\mathbb{R}^{|C|}$ be a generic sub–vector comprising the entries in ${\bf{z}}$ with indexes in $C\subset\{1,2,\ldots,m\}$, and denote with ${\bm{\xi}}_{C}\in\mathbb{R}^{|C|}$, $\bm{\Omega}_{CC}\in\mathbb{R}^{|C|\times|C|}$ and ${\bm{\Delta}}_{C\cdot}\in\mathbb{R}^{|C|\times q}$ the associated location sub–vector, dispersion sub–matrix and shape sub–matrix, respectively. Then
@@
The proof adapts the derivations in Arellano-Valle and Genton [10] to the parameterization considered in the present article. More specifically, by applying to (2) the linearity properties of elliptical distributions, it follows that
@@
thus obtaining the cumulative distribution function of the ${{\mbox{SUE}}}_{r,q}({\bf{A}}{\bm{\xi}}+\bm{b},{\bf{A}}\bm{\Omega}{\bf{A}}^{\top},{\bm{\Delta}}_{\bf{A}},\bm{\tau},\bar{\bm{\Gamma}},g^{(r+q)})$. The result for the marginals follows as a direct consequence by letting $\bm{b}=\bm{0}$ and setting ${\bf{A}}$ equal to a suitably–defined binary selection matrix such that ${\bf{A}}\bar{{\bf{z}}}={\bf{z}}_{C}$. ∎
@@
and conditional density generator given by $\smash{g^{(m_{i}+q)}_{Q_{j}({\bf{z}}_{j})}(u)}=g^{(m+q)}[Q_{j}({\bf{z}}_{j})+u]/g^{(m_{j})}[Q_{j}({\bf{z}}_{j})]$, for every $u\in\mathbb{R}$, where $Q_{j}({\bf{z}}_{j})$ is equal to $Q_{j}({\bf{z}}_{j})=({\bf{z}}_{j}-{\bm{\xi}}_{j})^{\top}\bm{\Omega}_{jj}^{-1}({\bf{z}}_{j}-{\bm{\xi}}_{j})$.
@@
$([\bar{{\bf{z}}}_{i}^{\top},(-\bm{\gamma}_{i\mid j}^{-1}\bar{{\bf{z}}}_{0})^{\top}]^{\top}\mid\bar{{\bf{z}}}_{j})\sim\mbox{{EC}}_{m_{i}+q}\left(\begin{bmatrix}{\bm{\xi}}_{i\mid j}\\ -\bm{\tau}_{i\mid j}\end{bmatrix},\begin{bmatrix}\bm{\Omega}_{i\mid j}&-\bm{\omega}_{i\mid j}{\bm{\Delta}}_{i\mid j}\\ -{\bm{\Delta}}_{i\mid j}^{\top}\bm{\omega}_{i\mid j}&\bar{\bm{\Gamma}}_{i\mid j}\end{bmatrix},g^{(m_{i}+q)}_{Q_{j}(\bar{{\bf{z}}}_{j})}\right),\quad\mbox{with}\ \ \ Q_{j}(\bar{{\bf{z}}}_{j})=(\bar{{\bf{z}}}_{j}-{\bm{\xi}}_{j})^{\top}\bm{\Omega}_{jj}^{-1}(\bar{{\bf{z}}}_{j}-{\bm{\xi}}_{j}),$
which also implies $\smash{(-\bm{\gamma}_{i\mid j}^{-1}\bar{{\bf{z}}}_{0}\mid\bar{{\bf{z}}}_{j})\sim\mbox{{EC}}_{q}(-\bm{\tau}_{i\mid j},\bar{\bm{\Gamma}}_{i\mid j},g^{(q)}_{Q_{j}(\bar{{\bf{z}}}_{j})})}$, as a direct consequence of the closure under marginalization. Combining these results with the selection representation in (2), and noticing that the event $-\bar{{\bf{z}}}_{0}\leq\bm{0}$ is equivalent to $-\bm{\gamma}_{i\mid j}^{-1}\bar{{\bf{z}}}_{0}\leq\bm{0}$ (since $\bm{\gamma}_{i\mid j}^{-1}$ is diagonal with positive entries), it follows that
@@
which coincides with the cumulative distribution function of the sue in (11) having parameters as in (12); see Equation (4) for the expression of the cumulative distribution function of a generic ${\mbox{SUE}}_{m,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},g^{(m+q)})$. Notice that the first equality in (14) follows from the fact that $\mathcal{p}({\bf{z}}_{i}\mid{\bf{z}}_{j})=\mathcal{p}({\bf{z}}_{i},{\bf{z}}_{j})/\mathcal{p}({\bf{z}}_{j})$, where $({\bf{z}}^{\top}_{i},{\bf{z}}^{\top}_{j})^{\top}={\bf{z}}$ is distributed as a sue and, by Lemma 1, the same holds for ${\bf{z}}_{j}$. Hence, from the selection representation in (2), it follows that
@@
Lemma 2 guarantees that when ${\bf{z}}=({\bf{z}}^{\top}_{i},{\bf{z}}^{\top}_{j})^{\top}$ is distributed as a sue, then also the conditional distribution for a generic sub–vector ${\bf{z}}_{i}$ belongs to the same family. Such a result conditions on a given realization ${\bf{z}}_{j}$ for the remaining entries in ${\bf{z}}$. In this respect, Lemma 3 states a novel finding which clarifies that the closure properties established in Lemma 2 can also be preserved when conditioning on a truncation event. A related result can be found in Proposition 4 of [6], but only with a focus on the sun sub–class. Here, such a property is proved for the whole sue family.

### v3

@@ 2.1 Multivariate unified skew-elliptical distributions
${\bf{z}}\stackrel{{\scriptstyle d}}{{=}}{\bm{\xi}}+\bm{\omega}{\bf{z}}^{\star},\qquad\qquad{\bf{z}}^{\star}\stackrel{{\scriptstyle d}}{{=}}(\tilde{\bf{z}}\mid\tilde{\bf{z}}_{0}+\bm{\tau}>\bm{0}),\qquad\begin{bmatrix}\tilde{\bf{z}}\\ \tilde{\bf{z}}_{0}\end{bmatrix}\sim\mbox{{EC}}_{m+q}\left(\begin{bmatrix}\bm{0}\\ \bm{0}\end{bmatrix},\begin{bmatrix}\bar{\bm{\Omega}}&{\bm{\Delta}}\\ {\bm{\Delta}}^{\top}&\bar{\bm{\Gamma}}\end{bmatrix},g^{(m+q)}\right).$ (5)
Such a representation is the one adopted by [5] for the sub-family of SUN distributions and is particularly convenient for deriving the mean vector and covariance matrix of ${\bf{z}}$. More specifically, leveraging (5), the law of total expectation, and the previously-discussed closure properties of elliptical distributions, we have that
@@
where $\psi$ is a scalar whose specific form can be obtained from the derivations in Arellano-Valle and Genton [10]. The above expressions clarify that moments of SUE random vectors can be directly obtained from those of multivariate truncated elliptical distributions [e.g., 13, 54, 39, 53, 59]. In addition, (7) shows that to enforce a lack of correlation among the entries in ${\bf{z}}$, it is not sufficient to impose suitable diagonal or block-diagonal structures within $\bm{\Omega}$. Rather, these constraints should be combined with additional ones, e.g., on the shape matrix ${\bm{\Delta}}$. Such a result is useful in Section 3 to derive examples of practically-meaningful priors and likelihoods inducing SUE posterior distributions. To this end, Sections 3.1, 3.2 and 3.3 state general conjugacy properties and then specialize these results to the two most widely-implemented examples of SUE distributions that are presented in detail in Section 2.2 below, namely, multivariate unified skew-normals (SUN) [5] and multivariate unified skew-$t$ (SUT) [62]. The results for SUN clarify that the conjugacy properties derived by Durante [32], Fasano and Durante [36] and Anceschi et al. [4] can be obtained as a special case, and under a different proof technique, of the more general SUE framework introduced in the present article. Conversely, the conjugacy results stated for SUT are a novel contribution that extends to a broader class of models of potential practical interest the findings in Song and Xia [57] and Zhang et al. [65] on specific Student’s $t$ linear regressions and multivariate probit formulations based on skew-elliptical link functions, respectively.
2.2 Relevant sub-classes of multivariate unified skew-elliptical distributions
@@
The SUN and SUT families arise from skewed perturbations of multivariate Gaussians and Student’s $t$ densities, respectively. As such, these formulations are arguably the most relevant and practically-impactful sub-classes in the SUE family. In addition, recalling Arellano-Valle and Azzalini [5], Arellano-Valle and Genton [10], and Wang et al. [62], both SUN and SUT admit additive stochastic representations which allow for i.i.d. sampling under posterior distributions belonging to these two sub-classes, thus facilitating Bayesian inference; see also Yin and Balakrishnan [63] for a recent extension of these stochastic representations to more general multivariate skew-elliptical distributions, beyond SUN and SUT. Sections 2.2.12.2.2 provide a concise overview of SUN and SUT sub-classes, respectively. A more extensive treatment can be found in, e.g., [5, 62].
2.2.1 Multivariate unified skew-normal distributions
@@
The SUN family has been introduced by Arellano-Valle and Azzalini [5] to provide a single class of distributions capable of unifying several extensions of the original multivariate skew-normal [19]. Relevant examples of representations that belong to such a wide class are extended multivariate skew-normals [15, 14] and closed skew-normals [42, 43], among others. Recalling Arellano-Valle and Azzalini [5], a random vector ${\bf{z}}\in\mathbb{R}^{m}$ has multivariate unified skew-normal (SUN) distribution, i.e., ${\bf{z}}\sim{\mbox{SUN}}_{m,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}})$, if its density $\mathcal{p}({\bf{z}})$ can be expressed as
@@
Although the above representation is originally derived in Arellano-Valle and Azzalini [5] under a selection representation similar to (5) applied to an underlying Gaussian vector, the density in (8) can also be derived directly from (1) under a suitable choice of the density generators. In particular, for $u\geq 0$, define $g^{(m)}=\phi^{(m)}(u)=(2\pi)^{-m/2}\exp(-u/2)$, $g^{(q)}=\phi^{(q)}(u)=(2\pi)^{-q/2}\exp(-u/2)$ and recall also that, in the particular Gaussian setting, the conditional generator $\phi^{(q)}_{Q({\bf{y}})}$ coincides with the unconditional one $\phi^{(q)}$. Then, recalling related derivations in Arellano-Valle and Genton [10], and replacing these density generators in (1), directly yields expression (8), thus clarifying that SUNs are special cases of SUE distributions.
Recalling the discussion in Section 1, the SUN family has been at the basis of important recent advancements in conjugate Bayesian inference for a broad class of routinely-implemented representations relying on fully-observed, discretized or partially-discretized Gaussian and multivariate skew-normal models [32, 36, 4]. As discussed in Section 3, these conjugacy properties can be recovered as a special case of those we derive for the general SUE family, which, in turn, allow us to extend such results to larger classes, including the SUT one introduced in Section 2.2.2.
2.2.2 Multivariate unified skew-$t$ distributions
The success of the SUN family has motivated several extensions aimed at deriving similar skewed representations for other sub-classes of the elliptical family. A noticeable and natural generalization is provided by the class of SUT distributions [e.g., 62] which can be obtained by replacing Gaussians with suitably-defined multivariate Student’s $t$ within the original selection representation of SUNs. Recalling, e.g., Wang et al. [62], this yields a density for the multivariate unified skew-$t$ (SUT) random vector ${\bf{z}}\sim{\mbox{SUT}}_{m,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},\nu)$, defined as
@@ 2.2.2 Multivariate unified skew-$t$ distributions
where $\alpha_{\nu,Q({\bf{z}})}=[\nu+Q({\bf{z}})]/{(\nu+m)}$, $Q({\bf{z}})=({\bf{z}}-{\bm{\xi}})^{\top}\bm{\Omega}^{-1}({\bf{z}}-{\bm{\xi}})$, and $\nu>0$ are the degrees of freedom. The remaining parameters have, likewise, similar roles and related interpretations to those in (1). Analogously, $t_{m}({\bf{z}}-{\bm{\xi}};\bm{\Omega},\nu)$, $T_{q}[{\alpha_{\nu,Q({\bf{z}})}^{-1/2}}\{\bm{\tau}+{\bm{\Delta}}^{\top}\bar{\bm{\Omega}}^{-1}\bm{\omega}^{-1}({\bf{z}}-{\bm{\xi}})\};\bar{\bm{\Gamma}}-{\bm{\Delta}}^{\top}\bar{\bm{\Omega}}^{-1}{\bm{\Delta}},\nu+m]$ and $T_{q}(\bm{\tau};\bar{\bm{\Gamma}},\nu)$ denote to the density and cumulative distribution functions, evaluated at the corresponding first argument, respectively, of the centered multivariate Student’s $t$ distributions with scale matrices $\bm{\Omega}\in\mathbb{R}^{m\times m}$, $\bar{\bm{\Gamma}}-{\bm{\Delta}}^{\top}\bar{\bm{\Omega}}^{-1}{\bm{\Delta}}\in\mathbb{R}^{q\times q}$ and $\bar{\bm{\Gamma}}\in\mathbb{R}^{q\times q}$, and degrees of freedom $\nu$, $\nu+m$ and $\nu$, respectively. Notice that, for $q=1$, one retrieves the multivariate extended skew-$t$ in [9]. Moreover, when $\bm{\tau}=\bm{0}$ and ${\bm{\Delta}}=\bm{0}$, the numerator and denominator in (9) coincide, and, therefore, the density reduces to that of a multivariate Student’s $t$ distribution $\mathcal{T}_{m}({\bm{\xi}},\bm{\Omega},\nu)$ with location ${\bm{\xi}}$, scale $\bm{\Omega}$, and degrees of freedom $\nu$. This implies that the multivariate Student’s $t$ is obtained as a special case of SUT.
@@
As for the SUN, also the SUT density within (9) can be derived from (1) under a suitable choice for the density generators. In particular, let $g^{(m)}=t_{v}^{(m)}(u)=c(\nu,m)[1+u/\nu]^{-(\nu+m)/2}$, and $g^{(q)}=t_{v}^{(q)}(u)=c(\nu,q)[1+u/\nu]^{-(\nu+q)/2}$, where the generic quantity $c(a,b)$ is defined as $c(a,b)=\Gamma[(a+b)/2]/[\Gamma(a/2)(\pi a)^{b/2}]$, while $\Gamma(\cdot)$ denotes the usual gamma function. Then, recalling, for example, Arellano-Valle and Genton [10], by replacing the generic density generators in (1) with those defined above and with the induced conditional density generator, yields, as a result of straightforward calculations, the SUT density in (9). Interestingly, as clarified in, e.g., Wang et al. [62], when $\nu\rightarrow\infty$, this density reduces to (8), thereby establishing a direct connection between SUT and SUN distributions. This suggests that the conjugacy properties derived in Durante [32], Fasano and Durante [36] and Anceschi et al. [4] for SUN might extend to SUT, and, more generally, to SUE distributions. Two promising results in this direction have been derived in Song and Xia [57] and in Zhang et al. [65], but only with a focus on Student’s $t$ linear regression and on specific multivariate binary data settings. Leveraging the SUE properties in Section 2.3, we prove in Section 3 that SUE conjugacy holds for a substantially larger class of models which further embraces formulations of potential interest in practice.
2.3 Properties of unified skew-elliptical distributions
@@
Lemmas 15 below state several central properties of SUE distributions that are at the core of the novel conjugacy results derived in Section 3. More specifically, Lemmas 12 establish closure under linear combinations, marginalization and conditioning of SUE distributions. All these properties have appeared also in Arellano-Valle and Genton [10], but under a different parameterization. Conversely, Lemmas 35 state novel results that are useful to study SUE conjugacy under broad classes of models and to derive special cases of potential interest in practical contexts.
@@
Let ${\bf{z}}\sim{{\mbox{SUE}}}_{m,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},g^{(m+q)})$. In addition, denote with ${\bf{A}}\in\mathbb{R}^{r\times m}$ a matrix with rank $r\leq m$, and let $\bm{b}\in\mathbb{R}^{r}$ be a vector of constants. Then
@@
where ${\bm{\Delta}}_{\bf{A}}=\bm{\omega}_{\bf{A}}^{-1}{\bf{A}}\bm{\omega}{\bm{\Delta}}$ with $\bm{\omega}_{\bf{A}}=\mbox{diag}({\bf{A}}\bm{\Omega}{\bf{A}}^{\top})^{1/2}$. Moreover, let ${\bf{z}}_{C}\in\mathbb{R}^{|C|}$ be a generic sub-vector comprising the entries in ${\bf{z}}$ with indexes in $C\subset\{1,\ldots,m\}$, and denote with ${\bm{\xi}}_{C}\in\mathbb{R}^{|C|}$, $\bm{\Omega}_{CC}\in\mathbb{R}^{|C|\times|C|}$ and ${\bm{\Delta}}_{C\cdot}\in\mathbb{R}^{|C|\times q}$ the associated location sub-vector, dispersion sub-matrix and shape sub-matrix, respectively. Then
@@
The proof adapts related derivations in Arellano-Valle and Genton [10] to the parameterization considered in the present article. More specifically, by applying to (2) the linearity properties of elliptical distributions, it follows that
@@
for any ${{\bf{z}}}^{*}\in\mathbb{R}^{r}$, thus obtaining the cumulative distribution function of the ${{\mbox{SUE}}}_{r,q}({\bf{A}}{\bm{\xi}}+\bm{b},{\bf{A}}\bm{\Omega}{\bf{A}}^{\top},{\bm{\Delta}}_{\bf{A}},\bm{\tau},\bar{\bm{\Gamma}},g^{(r+q)})$. The result for the marginals follows as a direct consequence by letting $\bm{b}=\bm{0}$ and setting ${\bf{A}}$ equal to a suitably-defined binary selection matrix such that ${\bf{A}}{\bf{z}}={\bf{z}}_{C}$. ∎
@@
and conditional density generator ${g^{(m_{i}+q)}_{Q_{j}({\bf{z}}_{j})}(u)}=g^{(m+q)}[Q_{j}({\bf{z}}_{j})+u]/g^{(m_{j})}[Q_{j}({\bf{z}}_{j})]$, with $Q_{j}({\bf{z}}_{j})=({\bf{z}}_{j}-{\bm{\xi}}_{j})^{\top}\bm{\Omega}_{jj}^{-1}({\bf{z}}_{j}-{\bm{\xi}}_{j})$.
@@
$([\bar{{\bf{z}}}_{i}^{\top},(-\bm{\gamma}_{i\mid j}^{-1}\bar{{\bf{z}}}_{0})^{\top}]^{\top}\mid\bar{{\bf{z}}}_{j})\sim\mbox{{EC}}_{m_{i}+q}\left(\begin{bmatrix}{\bm{\xi}}_{i\mid j}\\ -\bm{\tau}_{i\mid j}\end{bmatrix},\begin{bmatrix}\bm{\Omega}_{i\mid j}&-\bm{\omega}_{i\mid j}{\bm{\Delta}}_{i\mid j}\\ -{\bm{\Delta}}_{i\mid j}^{\top}\bm{\omega}_{i\mid j}&\bar{\bm{\Gamma}}_{i\mid j}\end{bmatrix},g^{(m_{i}+q)}_{Q_{j}(\bar{{\bf{z}}}_{j})}\right),\qquad Q_{j}(\bar{{\bf{z}}}_{j})=(\bar{{\bf{z}}}_{j}-{\bm{\xi}}_{j})^{\top}\bm{\Omega}_{jj}^{-1}(\bar{{\bf{z}}}_{j}-{\bm{\xi}}_{j}),$
which also implies ${(-\bm{\gamma}_{i\mid j}^{-1}\bar{{\bf{z}}}_{0}\mid\bar{{\bf{z}}}_{j})\sim\mbox{{EC}}_{q}(-\bm{\tau}_{i\mid j},\bar{\bm{\Gamma}}_{i\mid j},g^{(q)}_{Q_{j}(\bar{{\bf{z}}}_{j})})}$, as a direct consequence of the closure under marginalization. Combining these results with the selection representation in (2), and noticing that the event $-\bar{{\bf{z}}}_{0}\leq\bm{0}$ is equivalent to $-\bm{\gamma}_{i\mid j}^{-1}\bar{{\bf{z}}}_{0}\leq\bm{0}$ (since $\bm{\gamma}_{i\mid j}^{-1}$ is diagonal with positive entries), it follows that
@@
which coincides with the cumulative distribution function of the SUE in (11) having parameters as in (12); see (4) for the expression of the cumulative distribution function of a generic ${\mbox{SUE}}_{m,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},g^{(m+q)})$. Notice that the first equality in (14) follows from the fact that $\mathcal{p}({\bf{z}}_{i}\mid{\bf{z}}_{j})=\mathcal{p}({\bf{z}}_{i},{\bf{z}}_{j})/\mathcal{p}({\bf{z}}_{j})$, where $({\bf{z}}^{\top}_{i},{\bf{z}}^{\top}_{j})^{\top}={\bf{z}}$ is distributed as a SUE and, by Lemma 1, the same holds for ${\bf{z}}_{j}$. Hence, from the selection representation in (2), it follows that
@@
Lemma 2 guarantees that when ${\bf{z}}=({\bf{z}}^{\top}_{i},{\bf{z}}^{\top}_{j})^{\top}$ is distributed as a SUE, then also the conditional distribution for a generic sub-vector ${\bf{z}}_{i}$ belongs to the same family. Such a result conditions on a given realization ${\bf{z}}_{j}$ for the remaining entries in ${\bf{z}}$. In this respect, Lemma 3 states a novel finding which clarifies that the closure properties established in Lemma 2 can be preserved also when conditioning on a truncation event. Similar results can be found in [6] and [62], but with a focus on SUN and SUT sub-classes, respectively. Here, this property is proved for the whole SUE family.